OpenStudy (anonymous):

Determine the sum of a geometric series whose first term is 128, whose common ratio is ½ and whose number of terms is 7.

OpenStudy (anonymous):

you remember sum formula

OpenStudy (anonymous):

\[Sum _ {g.p} = \frac{a(1 - r^n)}{1 -r }\]

OpenStudy (anonymous):

\[Sum_{7} = \frac{128(1 - \frac{1}{2}^7)}{1 - \frac{1}{2}}\]

OpenStudy (anonymous):

254

OpenStudy (anonymous):

All you Need to do is to Calculate Good Luck : )

OpenStudy (anonymous):

Yeah it is 254 !

OpenStudy (anonymous):

THANKS MAN!!!

OpenStudy (anonymous):

Use the binomial coefficient formula to calculate the binomial coefficient for the 8th term in a polynomial expansion when the power of the binomial is 10

OpenStudy (anonymous):

sorry i forgot the formula : (